# Machine Translation Hallucination Detection for Low and High Resource Languages using Large Language Models

### v1

@@ Abstract
Recent advancements in massively multilingual machine translation systems have significantly enhanced translation accuracy; however, even the best performing systems still generate hallucinations, severely impacting user trust. Detecting hallucinations in Machine Translation (MT) remains a critical challenge, particularly since existing methods excel with High-Resource Languages (HRLs) but exhibit substantial limitations when applied to Low-Resource Languages (LRLs). This paper evaluates hallucination detection approaches using Large Language Models (LLMs) and semantic similarity within massively multilingual embeddings. Our study spans 16 language directions, covering HRLs, LRLs, with diverse scripts. We find that the choice of model is essential for performance. On average, for HRLs, Llama3-70B outperforms the previous state of the art by as much as 0.16 MCC (Matthews Correlation Coefficient). However, for LRLs we observe that Claude Sonnet outperforms other LLMs on average by 0.03 MCC. The key takeaway from our study is that LLMs can achieve performance comparable or even better than previously proposed models, despite not being explicitly trained for any machine translation task. However, their advantage is less significant for LRLs.
@@
Kenza Benkirane*1, Laura Gongas*1, Shahar Pelles1, Naomi Fuchs1, Joshua Darmon1, Pontus Stenetorp1, David Ifeoluwa Adelani1, Eduardo Sanchez1,2 1University College London (UCL), 2Meta *: Equal contribution
@@ 1 Introduction
Text generation models have drastically improved in recent years especially with the capabilities of LLMs in producing realistic and fluent output. However, hallucination continues to undermine user trust, as it generates and propagates misinformation and sometimes nonsensical outputs (Agarwal et al., 2018; Xu et al., 2023a; Guerreiro et al., 2023b).
@@
Previous work on post-translation evaluation has mainly focused on a single English-centric (EN) to a HRL direction, while studies including LRL remain limited (Raunak et al., 2022; Xu et al., 2023b). Recently, Dale et al. (2023) introduced HalOmi— a benchmark dataset for detecting hallucination in MT that includes ENHRLs (ten directions) and ENLRLs (six directions), as well as two non-English directions HRLLRL, including different scripts. BLASER-QE Communication et al. (2023), the state-of-the-art (SOTA) hallucination detector, is reported as the top performer on the HalOmi benchmark. It calculates a translation quality score by evaluating the similarity between encoded source texts and machine-translated texts within the SONAR embedding space Duquenne et al. (2023).
In this paper, we evaluate the performance of LLMs and embedding based methods as hallucination detectors, aiming to enhance performance in both HRLs and LRLs. To this end, we use the HalOmi benchmark dataset with a binary hallucination detection approach. For our evaluation, we include 14 methods: eight LLMs with different prompt variations, and four embedding spaces by computing the cosine similarity between source and translated texts.
@@ 2.1 Quality assessment of the dataset
The validation and test split was decided based on the translation direction. For the validation set, we selected the two translation directions DEEN, which encompasses 301 sentences. This choice was made as extensive resources and established benchmarks are available for this language pair Guerreiro et al. (2023a), with the expectation that the models would exhibit generalizability to less frequently used language pairs. For the test set, the other 16 pairs were used: more precisely, it includes four pairs with English and a HRL (ENAR , ENZH, ENRU, ENES), three pairs with English and a LRL (ENKS, ENMN, and ENYO), and one non-English HRL-LRL pair (ESYO). The test set includes 2,558 sentence pairs. This test set excludes six sentence pairs that were removed due to sensitive content flagged and filtered out by LLMs. A more detailed description of the dataset is available in Appendix B.
@@ 2.2 Hallucination detection setting
the classification of hallucinations was based on four severity levels: No Hallucination, Small Hallucination, Partial Hallucination, and Full Hallucination. This fine-grained categorization aimed to capture the nuances in the extent and impact of hallucinations on the translated output. We use this setting only as ablation study in Appendix C., both for consistency with the HalOmi benchmark, but also to assess the relevance of our binary detection approach.
@@
In this setting, all three instances of hallucinations were labelled as Hallucination, regardless of their severity. We also change the way the evaluation was done in HalOmi, with an appropriate prompt (Appendix D), and threshold calculation for binary classification for embeddings cosine similarity, see subsection 2.4. The primary reason for choosing this setting is the significant class imbalance in HalOmi, largely due to the scarcity of hallucinations across different severity levels. Some translation directions have particularly imbalanced data, for example ENRU, with the following distribution: out of 148 sentence pairs, we have 141 No Hallucination (96.6%), 1 Small (0.68%), 2 Partial (1.4%), and 4 Full (2.8%). High class imbalance can affect the ability of model to perform well (Prusa et al., 2016; Sordo and Zeng, 2005; Fernández et al., 2013).
@@ 2.3 LLMs for hallucination detection
We assessed the performances of eight LLMs, mixing capabilities models across LLMs families. We evaluate OpenAI’s GPT4-turbo and GPT4o; Cohere’s Command R and Command R+; Mistral’s Mistral-8x22b; Anthropic’s Claude Sonnet and Claude Opus and Meta’s Llama3-70B.111GPT3.5, Mistral Large and Llama3-8B were initially taken into account, but were excluded due to poor task understanding. More details about the selection are in subsection E.2.
@@
We derived the task description prompts from the Evaluate Hallucination and Evaluate Coherence in the Summarization Task prompts in G-Eval Liu et al. (2023). The CoT prompts were inspired by Evaluation Steps from G-Eval, and by the human annotation guidelines and severity level definitions from HalOmi. All prompts are available Appendix D. More details about the chosen hyperparameters with LLMs can be found in Appendix E.
@@ 3 Results
The results in Figure 2 demonstrate that LLMs have the best overall performance across languages for binary hallucination detection. Specifically, Llama3-70B surpasses the previous best performing model, BLASER-QE, by points, with an MCC of . For HRLs, 10 out of 12 evaluated methods outperform BLASER-QE (), with Llama3-70B greatly improving over the baseline by 16 points (). Notably, the results show that the choice of LLM should rely on the resource level; as for LRLs, Claude Sonnet achieves the highest average MCC. However, GPT4o was the more robust LLM across all languages, with the lowest standard deviation. Finally, for 13 out of the 16 evaluated translation directions, the evaluated methods outperform BLASER-QE, with the exception of KSEN, YOEN and ENMNI. Our findings on LLMs’ superior hallucination detection capabilities align with prior research on their effectiveness in MT quality assessment Kocmi and Federmann (2023).
@@
First, while Llama3-70B obtains the best performance overall, it was outperformed in most translation directions, especially in LRL. This result reveals a HRLs-centric approach of the model but also concludes that there is not one-LLM fits all resource levels. Secondly, for LRLs, models such as Sonnet, Opus, GPT4o, and Mistral —in order of decreasing performances, achieve higher scores, supporting the feasibility of employing LLMs in settings encompassing a wide range of languages. These results should be contrasted with a wide difference of hallucination distribution across resource levels, for example with the MNEN direction which only has No hallucination sentence pairs. More precisely, Sonnet and BLASER-QE perform on par for LRL, with the particularity that BLASER-QE has a significantly higher rate of false negatives, while Sonnet maintains a more balanced ratio of false positives to negatives. Moreover, BLASER-QE performs well in translations from English and comparably to Sonnet in translations to English, but falls short in non-English-centric translations, which follows the same trends as previously reported models in Dale et al. (2023). Figure 14 provides a more detailed view of these performance metrics.
@@ 4 Conclusion
In this work, we demonstrates that LLMs and embedding semantic similarity are highly effective for hallucination detection in machine translation, with LLMs establishing a new state-of-the-art performance across both high and low-resource languages. Our findings suggest that the optimal model selection depends on specific contexts, such as resource level, script, and translation direction. Our study highlight the need for further research to enhance hallucination detection capabilities, particularly in low-resource and non-English-centric translation settings.
@@
First, the dataset shows distribution imbalance across translation directions, with different trends for high and low resource languages, even after binarisation (see Appendix B): The HRLs show a pronounced data imbalance towards No Hallucination labels, with distribution between and . Moreover, for LRLs, there’s a broader interval, from to . This imbalance often results in models that classify translations as No hallucination being more frequently correct for HRLs than for LRLs, thereby introducing a bias into the binary evaluation. Moreover, the translation direction display a qualitative bias, as shown subsection B.3: HRLs and LRLs don’t have the same selection distribution which display a potential bias towards hallucination. Future dataset improvements should prioritize larger, more diverse samples, non-Latin scripts, and non-English centric translations. Using consistent source text across languages and balancing hallucination severity levels would enable more sophisticated methods, improve generalizability, and allow for a fair evaluation of models’ hallucination detection capabilities.
@@
The validation set used to identify the optimal threshold for non-LLM methods and the best prompt for LLMs only included ENDE translations. To improve parameter optimization and generalization across various translation directions, especially for low-resource languages (LRLs), cross-validation is recommended for future research, as suggested by Dale et al. (2023) and initially planned for our study. However, financial constraints associated with benchmarking non-open source models prevented the implementation of this approach. Future work should focus on developing novel approaches that perform well on well-studied high-resource languages (HRLs) while generalizing effectively to LRLs, assessing robustness, or exploring alternative methods to address this challenge within the limitations of dataset size.
@@
Finally, for benchmarking purposes, only the previous state-of-the-art (SOTA) was included for comparison against the newly evaluated methods. Therefore, for a more comprehensive analysis, it is recommended to include additional methods previously evaluated by HalOmi.
Appendix A Related Work
Significant advancements have been made in automatic machine translation evaluation, but these have predominantly focused on general translation errors. As a result, hallucinations are often overlooked, and evaluation scores may not reflect their impact due to their relatively low frequency compared to less severe errors like omissions Guerreiro et al. (2023a).
Previous studies have demonstrated that sentence similarity measures between source and translated texts, using cross-lingual embeddings such as LASER Heffernan et al. (2022) and LaBSE Feng et al. (2022), can effectively identify severe hallucinations Dale et al. (2022). However, the recently introduced Halomi dataset, A Manually Annotated Benchmark for Multilingual Hallucination and Omission Detection in Machine Translation Dale et al. (2023), which expands to include LRLs and non-English-centric translation directions, reveals the limitations of embedding semantic similarity methods primarily with LRLs. Conversely, the BLASER model Communication et al. (2023)—utilizing the SONAR embedding space Duquenne et al. (2023)—demonstrates greater robustness across language resources, establishing it as the latest state-of-the-art. This model notably improves performance in LRLs compared to previous methods, yet it still shows deficiencies in some non-English-centric directions.
Recent works have underlined the capabilities of LLMs in multilingual MT evaluation, demonstrating strong performances across various languages, although discrepancies are noted in LRLs Zhu et al. (2023); Xu et al. (2023b). G-Eval Liu et al. (2023) introduces a robust prompting framework for hallucination detection and demonstrates that LLMs can be used as automatic metrics to generate a single quality score. Furthermore, Kocmi and Federmann (2023) showed that LLMs, when appropriately prompted, can assess the quality of machine-generated translations, achieving state-of-the-art performance in system-level quality evaluation. Moreover, Fernandes et al. (2023) pioneered the evaluation of LLMs for MT tasks in LRLs using a new prompting technique, although its focus is primarily on broader translation errors rather than specifically on hallucination detection.
Appendix B Dataset description
B.1 Language acronyms mapping
@@
B.2 Hallucination distribution
B.2.1 Distribution of Hallucination in the severity ranking framework
@@
B.2.2 Distribution of Hallucination in the binary detection framework
@@ B.3 Selection distribution
Appendix C Ablation study
@@
C.1 LLMs for severity ranking
C.1.1 Prompt design
@@ C.1.2 Prompt evaluation
C.2 Embeddings for severity ranking
@@
C.3 Results
In the same way as in the binary detection setting, the validation results Table 4 allowed to select the otpimal prompt for each LLM, and then evaluate this best prompt across the test set, using here the ROC AUC score. Testing results aredisplayed Table 5, and presents ROC AUC scores for all methods per translation direction. For HRLs, embeddings’ high performance remains consistent with the binary hallucination approach. However, BLASER-QE remains the state-of-the-art in overall performance for severity ranking. The generalizability of these results requires further evaluation due to significant class imbalances in the dataset. Notably, in 11 of the 18 language directions, fewer than five samples are present in at least one hallucination severity category, see Appendix B.
@@ Appendix D Prompts
| Your task is to identify if the machine translated text has hallucination or not. Your task is to identify if the translated text is a hallucination. Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed. |
@@
Appendix E LLMs experiments
E.1 LLMs hyperparameters

### v3

@@ Abstract
Recent advancements in massively multilingual machine translation systems have significantly enhanced translation accuracy; however, even the best performing systems still generate hallucinations, severely impacting user trust. Detecting hallucinations in Machine Translation (MT) remains a critical challenge, particularly since existing methods excel with High-Resource Languages (HRLs) but exhibit substantial limitations when applied to Low-Resource Languages (LRLs). This paper evaluates sentence-level hallucination detection approaches using Large Language Models (LLMs) and semantic similarity within massively multilingual embeddings. Our study spans 16 language directions, covering HRLs, LRLs, with diverse scripts. We find that the choice of model is essential for performance. On average, for HRLs, Llama3-70B outperforms the previous state of the art by as much as 0.16 MCC (Matthews Correlation Coefficient). However, for LRLs we observe that Claude Sonnet outperforms other LLMs on average by 0.03 MCC. The key takeaway from our study is that LLMs can achieve performance comparable or even better than previously proposed models, despite not being explicitly trained for any machine translation task. However, their advantage is less significant for LRLs. 111Data and code are available on GitHub.
@@
Kenza Benkirane*1, Laura Gongas*1, Shahar Pelles1, Naomi Fuchs1, Joshua Darmon1, Pontus Stenetorp1, David Ifeoluwa Adelani2,3, Eduardo Sánchez1,4 1University College London, 2Mila - Quebec AI Institute , 3McGill University, 4Meta *: Equal contribution
@@ 1 Introduction
Text generation models have drastically improved in recent years especially with the capabilities of LLMs in producing realistic and fluent output. However, hallucination continues to undermine user trust, as it generates and propagates misinformation and sometimes nonsensical outputs (Agarwal et al., 2018; Xu et al., 2023a; Guerreiro et al., 2023b). This issue is especially critical in high-stakes domains like medicine and law, where hallucinations in medical texts can result in harmful misunderstandings about diagnoses or treatment instructions, and inaccuracies in legal contract translations can lead to severe financial or legal consequences.
@@
Previous work on post-translation evaluation has primarily focused on general translation errors, with evaluation scores often under-representing the impact of hallucinations due to their relatively low frequency compared to less severe errors like omissions Guerreiro et al. (2023a). Studies on hallucinations have mainly concentrated on English-centric (EN) to HRL translation direction, while research involving non-English LRLs remains limited (Raunak et al., 2022; Xu et al., 2023b).
For instance, sentence similarity measures between source and translated texts using cross-lingual embeddings, such as LASER Heffernan et al. (2022) and LaBSE Feng et al. (2022), have proven effective at identifying severe hallucinations Dale et al. (2022), though their limitations with LRLs are often overlooked Dale et al. (2023). Recent studies have highlighted the capabilities of LLMs in multilingual MT evaluation, demonstrating strong performances across various languages, although discrepancies remain for LRLs Zhu et al. (2023); Xu et al. (2023b). Kocmi and Federmann (2023) demonstrated that, when properly prompted, LLMs can assess the quality of machine-generated translations, achieving state-of-the-art results in system-level evaluation for HRLs. Furthermore, Fernandes et al. (2023) pioneered the use of LLMs for MT tasks in LRLs with the introduction of the AUTOMQM prompting technique, though this approach primarily targets broader translation errors rather than focusing on hallucination detection.
Recently, Dale et al. (2023) introduced HalOmi— a benchmark dataset for detecting hallucination in MT that includes ENHRLs (ten directions) and ENLRLs (six directions), as well as two non-English directions HRLLRL, including different scripts. BLASER-QE Communication et al. (2023), the state-of-the-art (SOTA) hallucination detector, is reported as the top performer on the HalOmi benchmark. It calculates a translation quality score by evaluating the similarity between encoded source texts and machine-translated texts within the SONAR embedding space Duquenne et al. (2023).
In this paper, we evaluate the performance of LLMs and embedding based methods as hallucination detectors, aiming to enhance performance in both HRLs and LRLs. To this end, we use the HalOmi benchmark dataset with a binary sentence-level hallucination detection approach. For our evaluation, we include 14 methods: eight LLMs with different prompt variations, and four embedding spaces by computing the cosine similarity between source and translated texts.
@@ 2.1 Quality assessment of the dataset
The validation and test split was decided based on the translation direction. For the validation set, we selected the two translation directions DEEN, which encompasses 301 sentences. This choice was made as extensive resources and established benchmarks are available for this language pair Guerreiro et al. (2023a), with the expectation that the models would exhibit generalizability to less frequently used language pairs. For the test set, the other 16 pairs were used: more precisely, it includes four pairs with English and a HRL (ENAR , ENZH, ENRU, ENES), three pairs with English and a LRL (ENKS, ENMN, and ENYO), and one non-English HRL-LRL pair (ESYO). The test set includes 2,558 sentence pairs. This test set excludes six sentence pairs that were removed due to sensitive content flagged and filtered out by LLMs. A more detailed description of the dataset is available in Appendix A.
@@ 2.2 Hallucination detection setting
the classification of hallucinations was based on four severity levels: No Hallucination, Small Hallucination, Partial Hallucination, and Full Hallucination. This fine-grained categorization aimed to capture the nuances in the extent and impact of hallucinations on the translated output. We use this setting only as ablation study in Appendix B., both for consistency with the HalOmi benchmark, but also to assess the relevance of our binary detection approach.
@@
In this setting, all three instances of hallucinations were labelled as Hallucination, regardless of their severity. We also change the way the evaluation was done in HalOmi, with an appropriate prompt (Appendix C), and threshold calculation for binary classification for embeddings cosine similarity, see subsection 2.4. The primary reason for choosing this setting is the significant class imbalance in HalOmi, largely due to the scarcity of hallucinations across different severity levels. Some translation directions have particularly imbalanced data, for example ENRU, with the following distribution: out of 148 sentence pairs, we have 141 No Hallucination (96.6%), 1 Small (0.68%), 2 Partial (1.4%), and 4 Full (2.8%). High class imbalance can affect the ability of model to perform well (Prusa et al., 2016; Sordo and Zeng, 2005; Fernández et al., 2013).
@@ 2.3 LLMs for hallucination detection
We assessed the performances of eight LLMs, mixing capabilities models across LLMs families. We evaluate OpenAI’s GPT4-turbo and GPT4o; Cohere’s Command R and Command R+; Mistral’s Mistral-8x22b; Anthropic’s Claude Sonnet and Claude Opus and Meta’s Llama3-70B.222GPT3.5, Mistral Large and Llama3-8B were initially taken into account, but were excluded due to poor task understanding. More details about the selection are in subsection D.2.
@@
We derived the task description prompts from the Evaluate Hallucination and Evaluate Coherence in the Summarization Task prompts in G-Eval Liu et al. (2023). The CoT prompts were inspired by Evaluation Steps from G-Eval, and by the human annotation guidelines and severity level definitions from HalOmi. All prompts are available Appendix C, with Figure 3 showing one of the prompts used for binary detection. More details about the chosen hyperparameters with LLMs can be found in Appendix D.
@@ 3 Results
The results in Figure 2 and Figure 15 demonstrate that LLMs have the best overall performance across languages for binary hallucination detection. Specifically, Llama3-70B surpasses the previous best performing model, BLASER-QE, by points, with an MCC of . For HRLs, 10 out of 12 evaluated methods outperform BLASER-QE (), with Llama3-70B greatly improving over the baseline by 16 points (). Notably, the results show that the choice of LLM should rely on the resource level; as for LRLs, Claude Sonnet achieves the highest average MCC. However, GPT4o was the more robust LLM across all languages, with the lowest standard deviation. Finally, for 13 out of the 16 evaluated translation directions, the evaluated methods outperform BLASER-QE, with the exception of KSEN, YOEN and ENMNI. Our findings on LLMs’ superior hallucination detection capabilities align with prior research on their effectiveness in MT quality assessment Kocmi and Federmann (2023).
@@
First, while Llama3-70B obtains the best performance overall, it was outperformed in most translation directions, especially in LRL. This result reveals a HRLs-centric approach of the model but also concludes that there is not one-LLM fits all resource levels. Secondly, for LRLs, models such as Sonnet, Opus, GPT4o, and Mistral —in order of decreasing performances, achieve higher scores, supporting the feasibility of employing LLMs in settings encompassing a wide range of languages. These results should be contrasted with a wide difference of hallucination distribution across resource levels, for example with the MNEN direction which only has No hallucination sentence pairs. More precisely, Sonnet and BLASER-QE perform on par for LRL, with the particularity that BLASER-QE has a significantly higher rate of false negatives, while Sonnet maintains a more balanced ratio of false positives to negatives. Moreover, BLASER-QE performs well in translations from English and comparably to Sonnet in translations to English, but falls short in non-English-centric translations, which follows the same trends as previously reported models in Dale et al. (2023). Figure 2 and Figure 15 provide a more detailed view of these performance metrics.
@@ 4 Conclusion
In this work, we demonstrates that LLMs and embedding semantic similarity are highly effective for hallucination detection in machine translation, with LLMs establishing a new state-of-the-art performance across both high and low-resource languages. Our findings suggest that the optimal model selection depends on specific contexts, such as resource level, script, and translation direction. Our study highlights the practical advantage of reference-free models like LLMs, which allow real-time hallucination detection without relying on external knowledge Su et al. (2024); Zhang et al. (2023). However, there remains a critical need for further research to improve hallucination detection, particularly in low-resource and non-English-centric translation settings.
@@
First, the dataset shows distribution imbalance across translation directions, with different trends for high and low resource languages, even after binarisation (see Appendix A): The HRLs show a pronounced data imbalance towards No Hallucination labels, with distribution between and . Moreover, for LRLs, there’s a broader interval, from to . This imbalance often results in models that classify translations as No hallucination being more frequently correct for HRLs than for LRLs, thereby introducing a bias into the binary evaluation. Moreover, the translation direction display a qualitative bias, as shown subsection A.3: HRLs and LRLs don’t have the same selection distribution which display a potential bias towards hallucination. Future dataset improvements should prioritize larger, more diverse samples, non-Latin scripts, and non-English centric translations. Using consistent source text across languages and balancing hallucination severity levels would enable more sophisticated methods, improve generalizability, and allow for a fair evaluation of models’ hallucination detection capabilities.
It is important to note there is a possibility of test set contamination, which is a common challenge in LLM research when the full training data is not publicly available. This issue primarily affects HRLs, where LLMs are predominantly trained; therefore, the impact on LRLs performance is expected to be minimal.
@@
The validation set used to identify the optimal threshold for non-LLM methods and the best prompt for LLMs only included ENDE translations. To improve parameter optimization and generalization across various translation directions, especially for LRLs, cross-validation is recommended for future research, as suggested by Dale et al. (2023) and initially planned for our study. However, financial constraints associated with benchmarking non-open source models prevented the implementation of this approach. Future work should focus on developing novel approaches that excel on well-studied HRLs while generalizing effectively to LRLs, assessing robustness, or exploring alternative methods to address this challenge within the limitations of dataset size.
@@
Finally, for benchmarking purposes, only the previous state-of-the-art was included for comparison against the newly evaluated methods. Therefore, for a more comprehensive analysis, it is recommended to include additional methods previously evaluated by HalOmi. Moreover, the benchmark can be further strengthened by identifying fine-grained hallucination spans to enhance interpretability.
Acknowledgements
We gratefully acknowledge OpenAI for providing API credits through their Researcher Access API program to Masakhane for evaluating GPT-4 LLMs. We also extend our thanks to Cohere for offering API credits for the use of their LLMs. David Adelani is supported by Canada CIFAR AI Chair program.
Appendix A Dataset description
A.1 Language acronyms mapping
@@
A.2 Hallucination distribution
A.2.1 Distribution of Hallucination in the severity ranking framework
@@
A.2.2 Distribution of Hallucination in the binary detection framework
@@ A.3 Selection distribution
Appendix B Ablation study
@@
B.1 LLMs for severity ranking
B.1.1 Prompt design
@@ B.1.2 Prompt evaluation
B.2 Embeddings for severity ranking
@@
B.3 Results
In the same way as in the binary detection setting, the validation results Table 4 allowed to select the otpimal prompt for each LLM, and then evaluate this best prompt across the test set, using here the ROC AUC score. Testing results aredisplayed Table 5, and presents ROC AUC scores for all methods per translation direction. For HRLs, embeddings’ high performance remains consistent with the binary hallucination approach. However, BLASER-QE remains the state-of-the-art in overall performance for severity ranking. The generalizability of these results requires further evaluation due to significant class imbalances in the dataset. Notably, in 11 of the 18 language directions, fewer than five samples are present in at least one hallucination severity category, see Appendix A.
@@ Appendix C Prompts
| Your task is to identify if the machine translated text has hallucination or not. Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed. |
@@
Appendix D LLMs experiments
D.1 LLMs hyperparameters